Group Selection Page HELP Controls Group Data Analysis Group ITDM WP76 DET Vacuum EEE WP78 WP74 FXE SPB MID HED SQS SXP Sample Environment Operation PSPO XO TS Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2025 logbook, Contact: F.Wolff-Fabris, A.Galler, H.Sinn, Page 1 of 4  Not logged in ELOG logo
ID Date Author Group Subject Text Attachments
  69   14 May 2025, 09:39 Jan GruenertPRCStatusDelayed start of SA1 (and SA3) beam delivery

The
beam delivery to the North branch (SASE1
/ SASE3) was delayed by 24h compared to the
 JG_2025-05-14_um_09.37.38.png 
  68   13 May 2025, 07:00 Raphael de WijnSPBExperiment summaryEnd of in-house beamtime

X-ray delivery


9.3 keV, 2.2-2.5 mJ pulse
  
  67   09 May 2025, 06:59 Raphael de WijnSPBShift summary
Shift
summary in-house night 2

X-ray delivery
  
  66   08 May 2025, 04:56 Raphael de WijnSPBShift summaryShift summary in-house beamtime night 1

X-ray delivery


9.3 keV, 1.7-2.2 mJ pulse
  
  65   05 May 2025, 22:41 Tokushi SatoSPBShift summaryX-ray
delivery


8.0 keV, 1300-1600 uJ pulse
  
  64   04 May 2025, 23:14 Tokushi SatoSPBShift summaryX-ray
delivery


8.0 keV, 1300 uJ pulse
  
  63   03 May 2025, 23:13 Tokushi SatoSPBShift summaryX-ray
delivery


8.0 keV, 1600 uJ pulse
  
  62   03 May 2025, 06:47 Tokushi SatoSPBShift summary**X-ray delivery**
- 8.3 , 1600 uJ pulse energy

**Optical laser delivery**
  
  61   02 May 2025, 08:07 Andreas KochXPDIssueConcerning eLog entry #59, SPB PopIn:

The problem of the motor position,
  
  60   01 May 2025, 23:35 Tokushi SatoSPBShift summaryX-ray
delivery


8.3 / 9.3 keV, 1400 uJ
  
  59   01 May 2025, 07:25 Romain LetrunSPBIssueSPB pop-in monitor was not dricing to the
switch when we retracted it to the "out
" position with the MDL this morning.
 Screenshot_from_2025-05-01_07-21-57.png 
  58   30 Apr 2025, 22:46 Romain LetrunSPBShift summaryX-ray delivery


9.3 keV, 1800 uJ pulse
energy

1 - 202 pulses at
  
  57   26 Apr 2025, 06:55 Tokushi SatoSPBShift summaryStaff: PV, TS, JK + external collaborators 

X-ray delivery:

10 keV, ~ 800 uJ, 128 pulses, tuning
  
  56   25 Apr 2025, 06:48 Tokushi SatoSPBShift summaryX-ray delivery:

10 keV, ~ 850 uJ, 128 pulses, tuning
was needed for 1 hour 
  
  55   24 Apr 2025, 08:42 Konopkova ZuzanaHEDIssueself-seeding level in SA2 is currently
very low (~100 uJ, about 50% in the pedestal),
need more information on how to proceed -
  
  54   24 Apr 2025, 06:56 Jayanath KoliyaduSPBShift summaryX-ray delivery:

10 keV, ~ 900 uJ, 128 pulses, tuning
was needed for few hours
  
  53   17 Apr 2025, 01:39 Chan KimSPBShift summaryX-ray delivery


18.8 keV, 260 uJ pulse
energy


Optical
  
  52   14 Apr 2025, 23:08 Tokushi SatoSPBShift summaryX-ray delivery:

20 keV,  250-400 uJ, 250 pulses,
stable delivery
  
  51   14 Apr 2025, 09:50 Jan GruenertXPDStatusBeam Downtime

During last night, apparently the
photon beam was down in all SASE beamlines
 JG_2025-04-14_um_09.43.03.png 
  50   13 Apr 2025, 22:55 Tokushi SatoSPBShift summaryStaff: PV, TS, JK, FK, + users

X-ray delivery:

20 keV, ~ 300-400 uJ, 250 pulses,
  
ELOG V3.1.4-7c3fd00